Prove that if A and B are matrices such that A is symmetric, then (BA−1 ) T (A−1BT ) −1 = In.


1
Expert's answer
2021-05-24T16:15:20-0400

To prove: "(BA^{-1})^T(A^{-1}B^T)^{-1}\n=I"


As A is symmetric matrix "A^T=A"


Taking LHS-


"(BA^{-1})^T(A^{-1}B^T)^{-1}"


"=(A^{-1})^TB^T(B^T)^{-1}(A^{-1})^{-1}"


"=(A^{-1})^TA^T ~~~~~~~~~~~~~~~~~~~( As A^T=A)"


"=I"